• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name moffit

    Log entry time 03:03:28 on October 5,2002

    Entry number 87343

    keyword=Left and Right Q1s reset

    The left and Right Q1s have been reset.
    Setpoint for Q1L = 1512.60222
    Setpoint for Q1R = 2085.30913

    Cryo Coord. called and said he has raised the lower limit for the fill valve on the Left Q1 by 10% but he did not increase the lower limit for the Right Q1.